OpenStudy (anonymous):

-x^2-5x-6... How to solve.

OpenStudy (precal):

you mean how to factor? If you solve it, it has to be equal to something.

OpenStudy (precal):

\[-x^2-5x-6\] \[-1(x^2+5x+6)\] -1(x+3)(x+2) soltuions would be x=-3 and x=-2

OpenStudy (precal):

by the way, I set it equal to zero

OpenStudy (anonymous):

-( x^2 + 5x + 6) = 0 - ( x +3) ( x+2) = 0 =>x = -3, x = -2